The volume, V, of a given mass of a gas varies directly as the temperature, T, and inversely as the pressure, if V=231cm^3 when T=42 and P=20kg/cm^2, what is the volume when T=30 and P=15kg/cm?
Answer by Theo(13342)

V = k*T/P where:

V = volume
T = temperature
P = pressure
k = constant of variation.

When:

V = 231 cm^3
T = 42 degrees
P = 20kg/cm^2,

The formula becomes:

231 = k*42/20

Solve for k to get:

k = 110

The constant of variation is equal to 110.

When T = 30 and P = 15 and k = 110, the formula becomes:

V = 30*110/15 = 220 cm^3

Here's a reference that explains how the variation formulas work.